1. Solid calcium carbonate is synthesized from calcium oxide powder and carbon dioxide gas.


2. The reaction between calcium carbide (kalburo, CaC2) and water forms acetylene (C2H2) and

calcium hydroxide, Ca (OH)2 .


3. The combustion of the mineral pyrite (FeS2) in oxygen gas produces another mineral hematite

(Fe2O3) and sulfur dioxide.


4. The oxidation of phosphine (PH3) to phosphorus pentoxide (P2O5) is given by the chemical

equation: PH3 + O2 → P2O5 + H2O


5. Thermite welding is an exothermic reaction,which produces molten iron that permanently

binds metals as it cools. The process does not require heat or electricity. Aluminum powder is

simply mixed with ferric oxide powder. The reaction is accompanied by a release of heat, and produces aluminum oxide and molten iron.


1
Expert's answer
2022-02-10T23:23:02-0500

1. Ca(OH)2 + CO2 = CaCO3 + H2O.


2. Zn + 2HCl = ZnCl2 + H2.


3. K2Cr2O7 + 2AgNO3 → Ag2Cr2O7 + 2KNO3